what is the numbers when the sum is 16 and product is -1472

Answer by Alan3354(69443)

what is the numbers when the sum is 16 and product is -1472
---------------
x + y = 16 --> y = 16 - xx*y = -1472
-----
x*(16-x) = -1472
x^2 - 16x - 1472 = 0
Solve for x.